OpenStudy (anonymous):

Given a circle with an 8" radius, find the area of the smaller segment whose chord is 8" long.

OpenStudy (anonymous):

Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius. -------------- Since all sides are 8", they form an equilateral triangle. The area of the sector is pi*r^2*60/360 Subtract the area of the triangle, that leaves the sector. --------------- Sector = pi*64*/6 = 32pi/3 sq inches Area of triangle = 4*8*cos(30) = 16sqrt(3) Area of segment = (32pi/3) - 16sqrt(3)